# Efficient Algorithms for Finite $\mathbb{Z}$-Algebras

### v3

@@ 5. Computing Primitive Idempotents
and $f\equiv h\mod\operatorname{Rad}(0)^{2^{k}}$. Now assume that $e\mod\operatorname{Rad}(0)$ is a primitive idempotent and that $f=e^{\prime}+e^{\prime\prime}$ can be written as the sum of two orthogonal idempotents. Then we have $e^{\prime}\in\operatorname{Rad}(0)$ or $e^{\prime\prime}\in\operatorname{Rad}(0)$, since $e$ is primitive. But $\operatorname{Rad}(0)$ consists only of nilpotent elements. Therefore $e^{\prime}$ or $e^{\prime\prime}$ has to be zero. ∎
@@
The number of iterations in Algorithm 5.1 necessary to lift the idempotents can be bounded as follows.
Proposition 5.3.
@@
    Let $p_{1}^{e_{1}},\dots,p_{s}^{e_{s}}$ be the elementary divisors of $R$. Then Algorithm 5.1 terminates after at most $\lceil\log_{2}(r+e_{1}+\cdots+e_{s})\rceil$ steps.
Proof.
@@
Part (b) follows immediately from (a), since the length of the torsion is given by the number of elementary divisors $p_{i}^{e_{i}}$ counted with multiplicity $e_{i}$. ∎
In order to compute the primitive idempotents of $R=P/I$, we can now use Algorithm 5.1 to lift the idempotents of $R/\operatorname{Rad}(0)$. For the task of finding the primitive idempotents of $R/\operatorname{Rad}(0)$, we consider the minimal associated primes of $I$. Let us recall the following remark from [20].
Remark 5.4.
@@
Definition 5.5.
Let $R$ be a finite $\mathbb{Z}$-algebra, and let $\operatorname{minAss}(R)$ be the set of minimal associated prime ideals of $R$. A maximal subset of $\operatorname{minAss}(R)$ such that all corresponding prime ideals are part of the same connected component of $\operatorname{Spec}(R)$ is called a connected component of $\operatorname{minAss}(R)$.
Since $R$ is a finite $\mathbb{Z}$-algebra, the associated primes of $R$ are either of height $n$ and do not contain a non-zero integer, or they are of height $n+1$ and hence maximal ideals. Now Algorithm 5.6 determines the connected components of $\operatorname{minAss}(R)$.
Algorithm 5.6.
(Computing the Connected Components of $\operatorname{minAss}(R)$)
@@
This is an algorithm which computes a set $M=\{C_{1},\dots,C_{\nu}\}$ such that $C_{1},\dots,C_{\nu}$ are the connected components of $\operatorname{minAss}(R)$. It is in ZPPIF.
Proof.
@@
Let us now show that the algorithm is in ZPPIF. The associated primes of $R$ can be computed in ZPPIF using Algorithm 4.2. Then only two types of computations remain. Namely, we need to decide whether the sum of two primes is equal to $\langle 1\rangle$ and whether one prime ideal is contained in another. Both of these tasks can be achieved in polynomial time by Proposition 2.9. ∎
@@
    Algorithm 4.2 yields the minimal associated primes $\mathfrak{m}_{1}=\langle\bar{x},\bar{y}-1,3\rangle$ and $\mathfrak{m}_{2}=\langle\bar{x},\bar{y}+1,2\rangle$ of height 3, as well as and the minimal associated primes $\mathfrak{p}_{1}=\langle\bar{x},\bar{y}\rangle$ and $\mathfrak{p}_{2}=\langle\bar{y},\bar{x}+5\rangle$ of height 2.
@@
    Thus the connected components of $\operatorname{minAss}(R)$ are $\{\{\mathfrak{m}_{1}\},\{\mathfrak{m}_{2}\},\{\mathfrak{p}_{1},\mathfrak{p}_{2}\}\}$.
From the connected components of $\operatorname{minAss}(R)$ we can now derive the primitive idempotents of $R$.
Algorithm 5.8.
@@
    Compute the connected components $C_{1},\dots,C_{\nu}$ of $\operatorname{minAss}(R)$ using Alg. 5.6.
@@
    Compute $J=\bigcap_{\mathfrak{p}\in\operatorname{minAss}(R)}\mathfrak{p}$.
@@
    Using Algorithm 5.1, lift the idempotents $q_{1},\dots,q_{\nu}$ of $R/J$ to idempotents of $R$ and return them.
Proof.
For a proof of the correctness of this algorithm, we again refer to Section 4 of [20]. Let us analyze the complexity of this algorithm. Step (1) can be performed in ZPPIF using Algorithm 5.6. The remaining steps can performed in polynomial time by Proposition 2.9, Algorithm 2.10, and Algorithm 5.1. The number of iterations necessary to perform Algorithm 5.1 has a polynomial bound by Proposition 5.3. ∎
Example 5.9.
Let us continue Example 5.7 and compute the primitive idempotents of the finite $\mathbb{Z}$-algebra $\mathbb{Z}[x,y]/I$ with $I=\left\langle x^{2}+5x,xy,y^{2}-y,6y\right\rangle$.
@@
    Using Algorithm 5.6, we already computed the connected components $\{\mathfrak{m}_{1}\}$, $\{\mathfrak{m}_{2}\}$, and $\{\mathfrak{p}_{1},\mathfrak{p}_{2}\}\}$ of $\operatorname{minAss}(R)$ in Example 5.7.
@@
    Using Proposition 2.9.e, we calculate $J=\mathfrak{p}_{1}\cap\mathfrak{p}_{2}\cap\mathfrak{m}_{1}\cap\mathfrak{m}_{2}=I$.
@@
    Using Proposition 2.9.e, we compute $\mathfrak{p}_{1}\cap\mathfrak{p}_{2}=\left\langle\bar{y},\bar{x}^{2}+5\bar{x}\right\rangle$.
@@
    We apply Algorithm 2.10 to the $\mathbb{Z}$-linear map $R/I\rightarrow R/(\mathfrak{p}_{1}\cap\mathfrak{p}_{2})\times R/\mathfrak{m}_{1}\times R/\mathfrak{m}_{2}$ and obtain the preimages $\bar{y}+1$, $3\bar{y}$, $-2\bar{y}$ of $e_{1},e_{2},e_{3}$.
@@ 6. Explicit $\mathbb{Z}$-Algebra Presentations and Strong Gröbner Bases
($\ast$) $I=\langle x_{i}x_{j}-\textstyle\sum\limits_{k=0}^{n}c_{ijk}x_{k},\;\textstyle\sum\limits_{k=0}^{n}a_{\ell k}g_{k}\mid i,j=1,\dots,n,\;\ell=1,\dots,m\rangle$
be the ideal in $P=\mathbb{Z}[x_{1},\dots,x_{n}]$ encoding the information of an explicitly given $\mathbb{Z}$-algebra $R=P/I$ as in Remark 2.1.
@@
In the first subsection we show that a presentation $R=P/I$ with $I$ as above allows us to compute a strong Gröbner basis of $I$ in polynomial time. In the second subsection we prove that, conversely, if $R=P/I$ and we know a strong Gröbner basis of $I$, then we can calculate a presentation as in Remark 2.1 in polynomial time.
@@ 6.1. Computing a Strong Gröbner Basis of an Explicitly Given $\mathbb{Z}$-Algebra
Finally, the list $L$ is updated such that its $\sigma$-smallest element is always the $\sigma$-smallest term greater than $t$ and not divisible by the leading monomial of some element of $G$. Since $P/(I_{1}\cap\cdots\cap I_{s})$ is a finite $\mathbb{Z}$-module, there exists for every $i\in\{1,\dots,n\}$ a number $\alpha_{i}\geq 1$ such that $x_{i}^{\alpha_{i}}\in\operatorname{LT_{\sigma}}(I_{1}\cap\cdots\cap I_{s})$. Hence only a finite number of terms can be added to the list $L$. This proves that the procedure terminates. ∎
@@
The following table shows how Algorithm 6.3 can be applied to compute a strong $\sigma$-Gröbner basis of $I\cap J$. The first five rows of this table correspond to the elements of $N$. The algorithm considers the terms $1,y,x,y^{2},xy,x^{2}$ in this order. Rows 6–11 in the table correspond to the representation vectors computed in step (4) of each iteration.
@@
If the $\mathbb{Z}$-algebras $P/I_{i}$ determined by the ideals $I_{i}$ are given as in Remark 2.1, then it is not necessary to compute their strong Gröbner bases separately, as the following corollary shows.
Corollary 6.5.
@@
Suppose that a finite $\mathbb{Z}$-algebra $R$ is explicitly given as in Remark 2.1, and let $I$ be the ideal in $P$ such that $R=P/I$ and $I$ is of the form described in ()normal-∗($\ast$)( ∗ ). Let $\sigma$ be a degree compatible term ordering on $\mathbb{T}^{n}$. Then Algorithm 6.3 computes a strong $\sigma$-Gröbner basis of $I$ in polynomial time.
Proof.
As mentioned in Remark 6.2.a, the representation vector in step (4) of an element $f\in P$ can be obtained by simplifying every product of indeterminates using the structure constants. The linear equation in step (6) can be solved in polynomial time by Remark 2.4. The claim then follows from the fact that the number of iterations is bounded by the number of generators of $R$. ∎
@@ 6.2. Computing an Explicit Representation
Now let $I\subseteq P$ be an ideal such that $R=P/I$ is a finite $\mathbb{Z}$-algebra. Given a strong Gröbner basis of $I$ with respect to some term ordering $\sigma$, our goal is to compute a representation of $R$ as in Remark 2.1. The first step is to find a suitable system of $\mathbb{Z}$-module generators of $R$.
Proposition 6.6.
@@
Proof.
It suffices to show that the $\mathbb{Z}$-submodule $Q=\sum_{t\in\mathcal{O}_{\sigma}}\mathbb{Z}(t+I)=\sum_{t\in\mathcal{O}_{\sigma}}\mathbb{Z}t+I$ of $P$ is equal to $P$. Suppose that $Q\subsetneq P$, and let $f\in P\setminus Q$ be a polynomial with minimal leading term. Then $\operatorname{LT_{\sigma}}(f)\in\mathcal{O}$ would imply $f-\operatorname{LC_{\sigma}}(f)\operatorname{LT_{\sigma}}(f)\in P\setminus Q$ which would contradict the minimality of $f$. Hence we have $\operatorname{LT_{\sigma}}(f)\in L$. This means there exist a polynomial $g\in I$ and a term $t\in\mathbb{T}^{n}$ such that $\operatorname{LC_{\sigma}}(g)=1$ and $\operatorname{LT_{\sigma}}(f)=t\operatorname{LT_{\sigma}}(g)$. But then $f-\operatorname{LC_{\sigma}}(f)tg$ has smaller leading term, again contradicting the minimality of $f$. ∎
@@
Let $I$ be an ideal in $P$, let $\sigma$ be a term ordering on $\mathbb{T}^{n}$, let $G$ be a minimal strong $\sigma$-Gröbner basis of $I$, and let $\{t_{1},\dots,t_{k}\}=\mathbb{T}^{n}\setminus L$, where $L$ equals $\{m\in\operatorname{LM_{\sigma}}(I)\mid\operatorname{LC_{\sigma}}(m)=1\}$ as in Proposition 6.6. Consider the following instructions.
@@
Proof.
By Proposition 6.6, the residue classes of the terms $t_{1},\dots,t_{k}$ generate the $\mathbb{Z}$-module $P/I$. Assume that $t_{1}>_{\sigma}t_{2}>_{\sigma}\cdots>_{\sigma}t_{k}$, and consider the $\mathbb{Z}$-module homomorphism
@@
Consequently, there is an element $d=(d_{1},\dots,d_{k})\in U$ with $d_{1}=\cdots=d_{i-1}=0$ and $\ell d_{i}=c_{i}$ for some $\ell\in\mathbb{Z}$. The tuple $c-\ell d$ corresponds to a polynomial whose leading term is smaller than $\operatorname{LT_{\sigma}}(f)$. This shows $c-\ell d\in U$, but then we get $c\in U$ in contradiction to our assumption. Hence we have the equality $\operatorname{Ker}(\varphi)=\langle U\rangle$ and $\varphi$ induces the inverse of the desired isomorphism. ∎
@@
    Apply Algorithm 6.7 to compute generators of a submodule $V\subseteq\mathbb{Z}^{k}$ such that $P/I\cong\mathbb{Z}^{k}/V$.
@@
Proof.
The residue classes of $t_{1},\dots,t_{k}$ generate $P/I$ as a $\mathbb{Z}$-module by Proposition 6.6. Algorithm 6.7 then correctly computes $V$ such that $P/I\cong\mathbb{Z}^{k}/V$. Finally, we check that $\operatorname{NF}_{\sigma,I}(t_{i}t_{j})$ is of the form given in step (4). Let $s$ be a term not contained in $\{t_{1},\dots,t_{k}\}$. Suppose that $\operatorname{NF}_{\sigma,I}(t_{i}t_{j})$ contains a monomial $ds$ with $d\in\mathbb{Z}$. Then we have $ds\notin\operatorname{LM_{\sigma}}(I)$, and in particular $s\notin L$, a contradiction. The time complexity of each step is clearly polynomial. ∎
In conclusion, we can see that an explicit representation of a finite $\mathbb{Z}$-algebra $R$ as in Remark 2.1 is equivalent to knowing a strong Gröbner basis of an ideal $I$ in $P=\mathbb{Z}[x_{1},\dots,x_{n}]$ such that $R=P/I$. Traditionally, many of the algorithms presented in this paper were executed using the calculation of a strong Gröbner basis. However, as there is no polynomial time bound for a suitable version of Buchberger’s Algorithm, the complexity bounds shown here would be impossible if $R$ were only given via $R=P/I$. As explicit representations of the type described in Remark 2.1 occur in many contexts (see for instance [20]), we hope that the algorithms and complexity bounds developed here may prove useful.

### v5

@@ 5. Computing Primitive Idempotents
and $f\equiv h\mod\operatorname{Rad}(0)^{2^{k}}$. Now assume that $e\mod\operatorname{Rad}(0)$ is a primitive idempotent and that $f=e^{\prime}+e^{\prime\prime}$ can be written as the sum of two orthogonal idempotents. Then we have $e^{\prime}\in\operatorname{Rad}(0)$ or $e^{\prime\prime}\in\operatorname{Rad}(0)$, since $e$ is primitive. But $\operatorname{Rad}(0)$ consists only of nilpotent elements. Therefore $e^{\prime}$ or $e^{\prime\prime}$ has to be zero.
@@
The number of iterations in Algorithm 15 necessary to lift the idempotents can be bounded as follows.
Proposition 16.
@@
    Let $p_{1}^{e_{1}},\dots,p_{s}^{e_{s}}$ be the elementary divisors of $R$. Then Algorithm 15 terminates after at most $\lceil\log_{2}(r+e_{1}+\cdots+e_{s})\rceil$ steps.
Proof 5.2.
@@
Part (b) follows immediately from (a), since the length of the torsion is given by the number of elementary divisors $p_{i}^{e_{i}}$ counted with multiplicity $e_{i}$.
In order to compute the primitive idempotents of $R=P/I$, we can now use Algorithm 15 to lift the idempotents of $R/\operatorname{Rad}(0)$. For the task of finding the primitive idempotents of $R/\operatorname{Rad}(0)$, we consider the minimal associated primes of $I$. Let us recall the following remark from [20].
Remark 17.
@@
{defi}
Let $R$ be a finite $\mathbb{Z}$-algebra, and let $\operatorname{minPrimes}(R)$ be the set of minimal associated prime ideals of $R$. A maximal subset of $\operatorname{minPrimes}(R)$ such that all corresponding prime ideals are part of the same connected component of $\operatorname{Spec}(R)$ is called a connected component of $\operatorname{minPrimes}(R)$.
Since $R$ is a finite $\mathbb{Z}$-algebra, the associated primes of $R$ are either of height $n$ and do not contain a non-zero integer, or they are of height $n+1$ and hence maximal ideals. Now Algorithm 18 determines the connected components of $\operatorname{minPrimes}(R)$.
Algorithm 18.
(Computing the Connected Components of $\operatorname{minPrimes}(R)$)
@@
This is an algorithm which computes a set $M=\{C_{1},\dots,C_{\nu}\}$ such that $C_{1},\dots,C_{\nu}$ are the connected components of $\operatorname{minPrimes}(R)$. It is in ZPPIF.
Proof 5.3.
@@
Let us now show that the algorithm is in ZPPIF. The associated primes of $R$ can be computed in ZPPIF using Algorithm 14. Then only two types of computations remain. Namely, we need to decide whether the sum of two primes is equal to $\langle 1\rangle$ and whether one prime ideal is contained in another. Both of these tasks can be achieved in polynomial time by Proposition 7.
@@
    Algorithm 14 yields the minimal associated primes $\mathfrak{m}_{1}=\langle\bar{x},\bar{y}-1,3\rangle$ and $\mathfrak{m}_{2}=\langle\bar{x},\bar{y}+1,2\rangle$ of height 3, as well as and the minimal associated primes $\mathfrak{p}_{1}=\langle\bar{x},\bar{y}\rangle$ and $\mathfrak{p}_{2}=\langle\bar{y},\bar{x}+5\rangle$ of height 2.
@@
    Thus the connected components of $\operatorname{minPrimes}(R)$ are $\{\{\mathfrak{m}_{1}\},\{\mathfrak{m}_{2}\},\{\mathfrak{p}_{1},\mathfrak{p}_{2}\}\}$.
From the connected components of $\operatorname{minPrimes}(R)$ we can now derive the primitive idempotents of $R$.
Algorithm 19.
@@
    Compute the connected components $C_{1},\dots,C_{\nu}$ of $\operatorname{minPrimes}(R)$ using Alg. 18.
@@
    Compute $J=\bigcap_{\mathfrak{p}\in\operatorname{minPrimes}(R)}\mathfrak{p}$.
@@
    Using Algorithm 15, lift the idempotents $q_{1},\dots,q_{\nu}$ of $R/J$ to idempotents of $R$ and return them.
Proof 5.4.
For a proof of the correctness of this algorithm, we again refer to Section 4 of [20]. Let us analyze the complexity of this algorithm. Step (1) can be performed in ZPPIF using Algorithm 18. The remaining steps can performed in polynomial time by Proposition 7, Algorithm 8, and Algorithm 15. The number of iterations necessary to perform Algorithm 15 has a polynomial bound by Proposition 16.
{exa}
Let us continue Example 5 and compute the primitive idempotents of the finite $\mathbb{Z}$-algebra $\mathbb{Z}[x,y]/I$ with $I=\left\langle x^{2}+5x,xy,y^{2}-y,6y\right\rangle$.
@@
    Using Algorithm 18, we already computed the connected components $\{\mathfrak{m}_{1}\}$, $\{\mathfrak{m}_{2}\}$, and $\{\mathfrak{p}_{1},\mathfrak{p}_{2}\}\}$ of $\operatorname{minPrimes}(R)$ in Example 5.
@@
    Using Proposition 7.e, we calculate $J=\mathfrak{p}_{1}\cap\mathfrak{p}_{2}\cap\mathfrak{m}_{1}\cap\mathfrak{m}_{2}=I$.
@@
    Using Proposition 7.e, we compute $\mathfrak{p}_{1}\cap\mathfrak{p}_{2}=\left\langle\bar{y},\bar{x}^{2}+5\bar{x}\right\rangle$.
@@
    We apply Algorithm 8 to the $\mathbb{Z}$-linear map $R/I\rightarrow R/(\mathfrak{p}_{1}\cap\mathfrak{p}_{2})\times R/\mathfrak{m}_{1}\times R/\mathfrak{m}_{2}$ and obtain the preimages $\bar{y}+1$, $3\bar{y}$, $-2\bar{y}$ of $e_{1},e_{2},e_{3}$.
@@ 6. Explicit $\mathbb{Z}$-Algebra Presentations and Strong Gröbner Bases
$I=\langle x_{i}x_{j}-\textstyle\sum\limits_{k=0}^{n}c_{ijk}x_{k},\;\textstyle\sum\limits_{k=0}^{n}a_{\ell k}g_{k}\mid i,j=1,\dots,n,\;\ell=1,\dots,m\rangle$ ($\ast$)
be the ideal in $P=\mathbb{Z}[x_{1},\dots,x_{n}]$ encoding the information of an explicitly given $\mathbb{Z}$-algebra $R=P/I$ as in Remark 1.
@@
In the first subsection we show that a presentation $R=P/I$ with $I$ as above allows us to compute a strong Gröbner basis of $I$ in polynomial time. In the second subsection we prove that, conversely, if $R=P/I$ and we know a strong Gröbner basis of $I$, then we can calculate a presentation as in Remark 1 in polynomial time.
@@ 6.1. Computing a Strong Gröbner Basis of an Explicitly Given $\mathbb{Z}$-Algebra
Finally, the list $L$ is updated such that its $\sigma$-smallest element is always the $\sigma$-smallest term greater than $t$ and not divisible by the leading monomial of some element of $G$. Since $P/(I_{1}\cap\cdots\cap I_{s})$ is a finite $\mathbb{Z}$-module, there exists for every $i\in\{1,\dots,n\}$ a number $\alpha_{i}\geq 1$ such that $x_{i}^{\alpha_{i}}\in\operatorname{LT_{\sigma}}(I_{1}\cap\cdots\cap I_{s})$. Hence only a finite number of terms can be added to the list $L$. This proves that the procedure terminates.
@@
The following table shows how Algorithm 21 can be applied to compute a strong $\sigma$-Gröbner basis of $I\cap J$. The first five rows of this table correspond to the elements of $N$. The algorithm considers the terms $1,y,x,y^{2},xy,x^{2}$ in this order. Rows 6–11 in the table correspond to the representation vectors computed in step (4) of each iteration.
@@
If the $\mathbb{Z}$-algebras $P/I_{i}$ determined by the ideals $I_{i}$ are given as in Remark 1, then it is not necessary to compute their strong Gröbner bases separately, as the following corollary shows.
Corollary 22.
@@
Suppose that a finite $\mathbb{Z}$-algebra $R$ is explicitly given as in Remark 1, and let $I$ be the ideal in $P$ such that $R=P/I$ and $I$ is of the form described in ()($\ast$)( ∗ ). Let $\sigma$ be a degree compatible term ordering on $\mathbb{T}^{n}$. Then Algorithm 21 computes a strong $\sigma$-Gröbner basis of $I$ in polynomial time.
Proof 6.2.
As mentioned in Remark 20.a, the representation vector in step (4) of an element $f\in P$ can be obtained by simplifying every product of indeterminates using the structure constants. The linear equation in step (6) can be solved in polynomial time by Remark 3. The claim then follows from the fact that the number of iterations is bounded by the number of generators of $R$.
@@ 6.2. Computing an Explicit Representation
Now let $I\subseteq P$ be an ideal such that $R=P/I$ is a finite $\mathbb{Z}$-algebra. Given a strong Gröbner basis of $I$ with respect to some term ordering $\sigma$, our goal is to compute a representation of $R$ as in Remark 1. The first step is to find a suitable system of $\mathbb{Z}$-module generators of $R$.
Proposition 23.
@@
Proof 6.3.
It suffices to show that the $\mathbb{Z}$-submodule $Q=\sum_{t\in\mathcal{O}_{\sigma}}\mathbb{Z}(t+I)=\sum_{t\in\mathcal{O}_{\sigma}}\mathbb{Z}t+I$ of $P$ is equal to $P$. Suppose that $Q\subsetneq P$, and let $f\in P\setminus Q$ be a polynomial with minimal leading term. Then $\operatorname{LT_{\sigma}}(f)\in\mathcal{O}$ would imply $f-\operatorname{LC_{\sigma}}(f)\operatorname{LT_{\sigma}}(f)\in P\setminus Q$ which would contradict the minimality of $f$. Hence we have $\operatorname{LT_{\sigma}}(f)\in L$. This means there exist a polynomial $g\in I$ and a term $t\in\mathbb{T}^{n}$ such that $\operatorname{LC_{\sigma}}(g)=1$ and $\operatorname{LT_{\sigma}}(f)=t\operatorname{LT_{\sigma}}(g)$. But then $f-\operatorname{LC_{\sigma}}(f)tg$ has smaller leading term, again contradicting the minimality of $f$.
@@
Let $I$ be an ideal in $P$, let $\sigma$ be a term ordering on $\mathbb{T}^{n}$, let $G$ be a minimal strong $\sigma$-Gröbner basis of $I$, and let $\{t_{1},\dots,t_{k}\}=\mathbb{T}^{n}\setminus L$, where $L$ equals $\{m\in\operatorname{LM_{\sigma}}(I)\mid\operatorname{LC_{\sigma}}(m)=1\}$ as in Proposition 23. Consider the following instructions.
@@
Proof 6.4.
By Proposition 23, the residue classes of the terms $t_{1},\dots,t_{k}$ generate the $\mathbb{Z}$-module $P/I$. Assume that $t_{1}>_{\sigma}t_{2}>_{\sigma}\cdots>_{\sigma}t_{k}$, and consider the $\mathbb{Z}$-module homomorphism
@@
Consequently, there is an element $d=(d_{1},\dots,d_{k})\in U$ with $d_{1}=\cdots=d_{i-1}=0$ and $\ell d_{i}=c_{i}$ for some $\ell\in\mathbb{Z}$. The tuple $c-\ell d$ corresponds to a polynomial whose leading term is smaller than $\operatorname{LT_{\sigma}}(f)$. This shows $c-\ell d\in U$, but then we get $c\in U$ in contradiction to our assumption. Hence we have the equality $\operatorname{Ker}(\varphi)=\langle U\rangle$ and $\varphi$ induces the inverse of the desired isomorphism.
@@
    Apply Algorithm 24 to compute generators of a submodule $V\subseteq\mathbb{Z}^{k}$ such that $P/I\cong\mathbb{Z}^{k}/V$.
@@
Proof 6.5.
The residue classes of $t_{1},\dots,t_{k}$ generate $P/I$ as a $\mathbb{Z}$-module by Proposition 23. Algorithm 24 then correctly computes $V$ such that $P/I\cong\mathbb{Z}^{k}/V$. Finally, we check that $\operatorname{NF}_{\sigma,I}(t_{i}t_{j})$ is of the form given in step (4). Let $s$ be a term not contained in $\{t_{1},\dots,t_{k}\}$. Suppose that $\operatorname{NF}_{\sigma,I}(t_{i}t_{j})$ contains a monomial $ds$ with $d\in\mathbb{Z}$. Then we have $ds\notin\operatorname{LM_{\sigma}}(I)$, and in particular $s\notin L$, a contradiction. The time complexity of each step is clearly polynomial.
In conclusion, we can see that an explicit representation of a finite $\mathbb{Z}$-algebra $R$ as in Remark 1 is equivalent to knowing a strong Gröbner basis of an ideal $I$ in $P=\mathbb{Z}[x_{1},\dots,x_{n}]$ such that $R=P/I$. Traditionally, many of the algorithms presented in this paper were executed using the calculation of a strong Gröbner basis. However, as there is no polynomial time bound for a suitable version of Buchberger’s Algorithm, the complexity bounds shown here would be impossible if $R$ were only given via $R=P/I$. As explicit representations of the type described in Remark 1 occur in many contexts (see for instance [20]), we hope that the algorithms and complexity bounds developed here may prove useful.